# RBI-ThPhys-2024-15

### v1

@@ IV Simulation of di-Higgs events
To investigate interference effects between resonant and non-resonant contributions to pair production of SM-like Higgs bosons, we simulate events of di-Higgs production via gluon fusion ($gg\to hh$) both at leading-order (LO) and next-to-LO (NLO) in perturbative quantum chromodynamics (QCD). We perform the LO simulations with the MadGraph5_amc@nlo Alwall:2014hca event generator, using an implementation of the Two Real Singlet Extension of the SM Papaefstathiou:2020lyp with the model file from gitlabandreas . While this model contains three neutral Higgs bosons, we decouple one of these in our implementation. We denote the two non-decoupled scalar fields as $h$ and $H$, where $h$ is identified as the observed Higgs boson with a mass of $m_{h}=125\,{\,\rm GeV}$, and $H$ denotes the additional heavier Higgs boson with a variable mass $m_{H}$. The model leads to contributions to the di-Higgs amplitude from both non-resonant and resonant Feynman diagrams, including interference terms. Among the non-resonant diagrams, the box diagram, $\Box$, (Fig. 2, upper-left), is independent of the Higgs trilinear couplings, $\lambda_{i}$, whereas the $s$-channel diagram, $\mathrm{S_{h}}$, depends on the $\lambda_{hhh}$ coupling (Fig. 2, upper-right). The resonant $s$-channel contribution, $\mathrm{S_{H}}$, depends on the $\lambda_{Hhh}$ trilinear coupling (Fig. 2, bottom).
[…]
@@
The NLO simulations are performed using the POWHEG 2.0 Nason:2004rx ; Frixione:2007vw ; Alioli:2010xd event generator. Currently there is no POWHEG model that simulates di-Higgs events including both non-resonant and resonant diagrams. Instead we simulate these contributions separately neglecting the interference between resonant and non-resonant diagrams, and in Sect. VI we will employ a reweighting method to account for the missing interference terms. The non-resonant simulations follow the implementation described in Refs. Heinrich:2017kxx ; Heinrich:2019bkc ; Heinrich:2020ckp . The resonant simulation is performed by generating single $H$ events using the implementation presented in POWHEG Bagnaschi:2011tu , followed by the $H\rightarrow hh$ decay implemented in PYTHIA 8.310 Bierlich:2022pfr .
@@
    The $\phi$ and $\eta$ direction of the $b$-quarks are smeared independently by shifting the truth-level $\phi$ and $\eta$ on an case-by-case basis by an amount determined from randomly sampling a Gaussian function with a mean of 0 and and an uncertainty of $\sigma_{\phi}=\sigma_{\eta}=0.05$, which is the typical angular resolution for $b$-jets with $p_{\mathrm{T}}\approx 100\,{\,\rm GeV}$ estimated using simulations of the CMS detector.
@@
    Finally, the energy/momentum of each $b$-quark is smeared by scaling its 4-vector by $y$, where $y$ is determined by randomly sampling a Gaussian function with a mean of 1 and a width of $\sigma=0.15$. The 15% value is chosen to ensure that the overall resolution on the reconstructed $h$ mass, $m_{h}$, matches the resolution quoted by the CMS Collaboration in Ref. CMS:2019uxx .
We will refer to the smeared $b$-quarks as “$b$-jets” in the following. The Higgs boson candidates $h$ are reconstructed by pairing the $b$-jets. There are three possible combination of pairs, and we choose the combination that minimises $(m_{h_{1}}-125\,{\,\rm GeV})^{2}+(m_{h_{2}}-125\,{\,\rm GeV})^{2}$, where $m_{h_{1}}$ and $m_{h_{2}}$ are the reconstructed masses of the two $h$ candidates. We estimate the di-Higgs invariant mass, $m_{hh}$, following the method in Ref. ATLAS:2022hwc . This method improves the $m_{hh}$ resolution by scaling the four-momenta of each of the reconstructed $h$ such that $m_{h}=125\,{\,\rm GeV}$. Following this method, for signal events with $m_{H}=600\,{\,\rm GeV}$, the $m_{hh}$ resolution is about 6%.
@@ V Numerical investigation of interference effects
In this section, we analyze the parameter dependence of the LO QCD simulations performed with MadGraph5_amc@nlo. In the first part, we discuss in detail different observables and their interference behavior at two example parameter points. The second part contains a parameter scan, where we are mostly interested in the effect of the interference on the cross section.
To evaluate the impact of the interference term, we split the diagrams describing the total process $P_{\text{tot}}=pp\rightarrow hh$ into two type of subprocesses, defined as
@@
    $P_{\mathrm{S_{H}}}=pp\rightarrow H\rightarrow hh$
@@
To generate the $P_{\text{woH}}$ subprocesses, which include the $\Box$ and $\mathrm{S_{h}}$ diagrams as shown in Fig. 2, we set the coupling constant $\lambda_{Hhh}$ to zero in the simulation. For the widths of the Higgs bosons $H$ and $h$, we use an estimated tree-level total width, consisting of the partial widths for $h\rightarrow VV^{\left(*\right)}$, $h\rightarrow ff$, and $H\rightarrow hh$ and an effective $h\rightarrow\gamma\gamma$ partial width [^5] The total width as the sum of these partial widths is calculated using the AutoWidth feature of MadGraph5_amc@nlo.. As the current model file does not include the gluonic decay mode automatically, we also added this decay width by hand, where we use the full top quark mass dependence and perform the calculation at leading order. All partial widths were calculated with MadGraph5_amc@nlo. Some example values are shown in Tab. 1.
@@ V.1 Illustration of interference Effects
For illustrative purposes, we show here the results for two parameter points, namely the points $m_{H}=300\,{\,\rm GeV}$, $\sin\alpha$ =0.17 and $\tan\beta$ =3.3 and $m_{H}=600\,{\,\rm GeV}$, $\sin\alpha$ =0.17 and $\tan\beta$ =1.6. For these points, we simulated $10^{6}$ events for each subprocess, $P_{\text{tot}}$, $P_{\mathrm{S_{H}}}$ and $P_{\text{woH}}$. We investigate the interference effects for different observables directly by evaluating the difference between the coherent sum $P_{\text{tot}}$ of all contributions to $gg\to hh$ and the incoherent sum $P_{\mathrm{S_{H}}}+P_{\text{woH}}$. In particular, we analyze the $m_{hh}$, $p_{\mathrm{T}}$ and $\Delta\phi_{hh}$ distributions, where $\Delta\phi_{hh}$ is the angle between the final state Higgs bosons. Here, $p_{\mathrm{T}}$ is shown for both final state Higgs bosons. While the behaviour of the di-Higgs invariant mass is well known, and the corresponding interference effects have been discussed in detail in the literature (see e.g. DiMicco:2019ngk and references therein), other variables have not been investigated yet. However, all of these can in principle serve as input for both cut-based as well as neural-network based analyses. Therefore, it is imminent to equally be aware of interference effects in as many as possible differential distributions.
@@
As we will discuss in the following, $m_{hh}$ and $p_{\mathrm{T}}$ show interesting interference effects, as can be seen in Fig. 3, which shows the number of total events expected using full Run 2 luminosity = $\int\mathcal{L}\,=\,139\,{\rm fb}^{-1}$. In contrast, $\Delta\phi_{hh}$ is not significantly affected for these parameters points, and we chose to defer the corresponding discussion to Appendix A. We observe the well-known peak-dip structure around $m_{hh}\,\sim\,m_{H}$, that has already been discussed in the literature: for $m_{hh}$ slightly larger than the mass of the heavy scalar, there is a negative interference that leads to a smaller number of events than expected from incoherently adding the different distributions. A similar effect appears for $m_{hh}$ slightly below $m_{H}$ although the interference effect has a different sign there. These are therefore regions in the $m_{hh}$ distributions where interference effects can become important. The interference behaves similarly for the $p_{\mathrm{T}}$ distributions, with positive interference for $p_{\mathrm{T}}$ lower than the peak and negative interference for $p_{\mathrm{T}}$ higher than the $p_{\mathrm{T}}$ peak.
@@
Figure 3: The upper plots show the invariant mass $m_{hh}$ and the lower plots show the transversal momenta $p_{\mathrm{T}}$ of both final state Higgs bosons for the two parameter points. The blue line corresponds to $N_{\text{tot}}$, the red line corresponds to $N_{\mathrm{S_{H}}}$, the orange line corresponds to $N_{\text{woH}}$ and the green line is the sum $N_{\mathrm{S_{H}}}+N_{\text{woH}}$ neglecting interference effects. The event numbers $N_{i}$ corresponding to process $P_{i}$ were simulated assuming full Run 2 luminosity. Bin size is $10\,{\,\rm GeV}$.
@@
Another important question is whether these effects are statistically significant. In Fig. 4, we therefore show the relative difference for the event numbers for the full Run 2 luminosity. As a simple estimate, we only take statistical errors into account. We see that the interference effects are non-negligible over large regions of parameter space, and exceed the statistical uncertainty estimate we applied.
@@
Figure 4: These plots show the interference calculated from the difference between the blue line and the green line seen in Fig.  3 as the green line and a statistical error calculated as the $\sqrt{N_{\text{tot}}}$ of the $P_{\text{tot}}$ process assuming full Run 2 luminosity. The upper plots show this for the invariant mass $m_{hh}$ and the lower plots show this for the transversal momentum $p_{\mathrm{T}}$ for the two parameter points.
@@ V.2 Parameter scan
To investigate the effects of interference between the resonant and non-resonant diagrams, we examine the cross-section corresponding to the interference term for a broad range of parameters. For this purpose, we perform a parameter scan for a range of the relevant parameters in the parameter volume of $[260\,{\,\rm GeV},800\,{\,\rm GeV}]\times[-0.24,0.24]\times[0.5,4.0]$ of the $\{{m_{H},\sin\alpha,\tan\beta\}}$ space where we scan $m_{H}$ in steps of $60$ GeV, $\sin\alpha$ is steps of $0.08$ and $\tan\beta$ in steps of $0.5$. The resulting parameter points are displayed in Tab. 2. We omitted the points with $\sin\alpha=0$ as they represent the SM case where the second Higgs decouples, leading to no interference.
@@
Figure 5: The different bins on the x-axis show the investigated values of $m_{H}$. The points inside the bins are ordered by the corresponding $\tan\beta$ values from left to right. The y-axis shows $\sin\alpha$, where the SM case of $\sin\alpha=0$ is omitted. Circles represent allowed points and triangles represent the points that do not pass the scan constraints described in Sec. II. The maximal relative interference for allowed points is around $13\%$.
We investigated the interference effects for events with invariant mass $m_{hh}$ above or below $m_{H}$. For this we used the pylhe pylhe python package to divide our generated events into those with higher and those with lower invariant mass than the respective $m_{H}$. Accordingly, we split the cross sections into the interference for the lower invariant mass range, $\left(\Delta\sigma\right)_{\text{abs}}^{<}$, and for the higher invariant mass range, $\left(\Delta\sigma\right)_{\text{abs}}^{>}$, respectively, as well as their absolute sum
@@
and the corresponding relative interferences $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, $\left(\Delta\sigma\right)_{\text{rel}}^{>}$ and $\left(\Delta\sigma\right)_{\text{rel}}^{\sum}$, defined accordingly by normalization to the total cross section over the whole mass range, equivalent to Eq. (24). The results are displayed in Fig. 6f. As shown in Figs. 3a and especially  3b, the interference seems to have a positive value for $m_{hh}<m_{H}$ and a negative value for $m_{hh}>m_{H}$. Therefore, although the total interference integrated over the whole mass range might be small in some scenarios, effects in differential distributions could still be large. We therefore also provide the corresponding information on the different mass regions discussed above as well as the absolute value of deviations between naive summing of the different contributions and the total process including interference. Fig. 6a shows that a higher positive mixing angle enhances both $\left(\Delta\sigma\right)_{\text{abs}}^{<}$ and $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, see Fig. 6b. Furthermore, Fig. 6c displays that $\left(\Delta\sigma\right)_{\text{abs}}^{>}$ has the largest negative values for high mixing angles in the $m_{H}$ region around $380\,{\,\rm GeV}$, see also Fig. 6d. Fig. 6e and especially Fig. 6f show points, which appeared to have little interference based on Figs. 5a and 5b actually have two interference effects of opposite sign. In Sect. VII.2 we will introduce benchmark points that exhibit this feature (benchmarks BM2 and BM3).
@@
Figure 6: Interferences for invariant di-Higgs masses $m_{hh}$ above and below the heavy scalar mass $m_{H}$, separately. The bins on the x-axis show the different values of $m_{H}$ investigated and inside the bins the points are ordered by the corresponding $\tan\beta$ values from left to right. The $\sin\alpha$ value is shown on the y-axis, where $\sin\alpha=0$ is left out, as it corresponds to the SM case. Circles represent allowed points and triangles represent the points that do not pass the scan constraints described in Sect. II
@@
Figure 7: The different bins on the x-axis show the different values of $m_{H}$ investigated and inside the bins the points are ordered by the corresponding $\tan\beta$ values from left to right. The $\sin\alpha$ value is shown on the y-axis, where $\sin\alpha=0$ is left out, as it corresponds to the SM-case. Circles represent allowed points and triangles represent the points that do not pass the scan constraints described in Sect. II.
@@
. This in principle also manifests itself in discrepancies of the branching ratios of the $125\,{\,\rm GeV}$ scalar. However, as in our scenario the $125\,{\,\rm GeV}$ particle only appears in highly off-shell scenarios, we did not modify the total width but keep the LO calculated value from MadGraph in our simulations.

### v2

@@ IV Simulation of di-Higgs events
To investigate interference effects between resonant and non-resonant contributions to pair production of SM-like Higgs bosons, we simulate events of gluon fusion di-Higgs production ($gg\to hh$) both at LO and NLO in perturbative quantum chromodynamics (QCD). We perform the LO simulations with the MadGraph5_amc@nlo Alwall:2014hca event generator, using an implementation of the Two Real Singlet Extension of the SM Papaefstathiou:2020lyp with the model file from gitlabandreas . While this model contains three neutral Higgs bosons, we decouple one of these in our implementation. We denote the two non-decoupled scalar fields as $h$ and $H$, where $h$ is identified as the observed Higgs boson with a mass of $m_{h}=125\,{\,\rm GeV}$, and $H$ denotes the additional heavier Higgs boson with a variable mass $m_{H}$. The model leads to contributions to the di-Higgs amplitude from both non-resonant and resonant Feynman diagrams, including interference terms. Among the non-resonant diagrams, the box diagram, $\Box$, (Figure 2, upper-left), is independent of the Higgs trilinear couplings, $\lambda_{i}$, whereas the $s$-channel diagram, $\mathrm{S_{h}}$, depends on the $\lambda_{hhh}$ coupling (Figure 2, upper-right). The resonant $s$-channel contribution, $\mathrm{S_{H}}$, depends on the $\lambda_{Hhh}$ trilinear coupling (Figure 2, bottom).
Refer to caption           Refer to caption
@@
The NLO simulations are performed using the POWHEG 2.0 Nason:2004rx ; Frixione:2007vw ; Alioli:2010xd event generator. Currently there is no POWHEG model that simulates di-Higgs events including both non-resonant and resonant diagrams. Instead we simulate these contributions separately neglecting the interference between resonant and non-resonant diagrams, and in Section VI we will employ a reweighting method to account for the missing interference terms. The non-resonant simulations follow the implementation described in Refs. Heinrich:2017kxx ; Heinrich:2019bkc ; Heinrich:2020ckp . The resonant simulation is performed by generating single $H$ events using the implementation presented in POWHEG Bagnaschi:2011tu , followed by the $H\rightarrow hh$ decay implemented in PYTHIA 8.310 Bierlich:2022pfr .
@@
    The $\phi$ and $\eta$ direction of the $b$-quarks are smeared independently by shifting the truth-level $\phi$ and $\eta$ on an case-by-case basis by an amount determined from randomly sampling a Gaussian function with a mean of 0 and and an uncertainty of $\sigma_{\phi}=\sigma_{\eta}=0.05$, which is the typical angular resolution for $b$-jets with $p_{\mathrm{T}}\approx 100\,{\,\rm GeV}$ as estimated using simulations of the CMS detector.
@@
    Finally, the energy/momentum of each $b$-quark is smeared by scaling its 4-vector by $y$, where $y$ is determined by randomly sampling a Gaussian function with a mean of 1 and a width of $\sigma=0.15$. The 15% value is chosen to ensure that the overall resolution on the reconstructed $h$ mass matches the resolution quoted by the CMS Collaboration in Ref. CMS:2019uxx .
We will refer to the smeared $b$-quarks as “$b$-jets” in the following. The $h$ candidates are reconstructed by pairing the $b$-jets. There are three possible combination of pairs, and we choose the combination that minimises $(m_{h_{1}}-125\,{\,\rm GeV})^{2}+(m_{h_{2}}-125\,{\,\rm GeV})^{2}$, where $m_{h_{1}}$ and $m_{h_{2}}$ are the reconstructed masses of the two $h$ candidates. We estimate the di-Higgs invariant mass, $m_{hh}$, following the method in Ref. ATLAS:2022hwc . This method improves the $m_{hh}$ resolution by scaling the four-momenta of each of the reconstructed $h$ such that $m_{h}=125\,{\,\rm GeV}$. Following this method, for signal events with $m_{H}=600\,{\,\rm GeV}$, the $m_{hh}$ resolution is about 6%.
@@ V Numerical investigation of interference effects
In this section, we analyse the parameter dependence of the LO QCD simulations performed with MadGraph5_amc@nlo. In the first part, we discuss in detail different observables and their interference behaviour at two example parameter points. The second part contains a parameter scan, where we are mostly interested in the effect of the interference on the cross-section.
To evaluate the impact of the interference term, we split the diagrams describing the total process $P_{\text{tot}}=pp\rightarrow hh$ into two subprocesses, defined as
@@
    $P_{\mathrm{S_{H}}}=pp\rightarrow H\rightarrow hh$,
@@
To generate the $P_{\text{woH}}$ subprocesses, which include the $\Box$ and $\mathrm{S_{h}}$ diagrams as shown in Figure 2, we set the coupling constant $\lambda_{Hhh}$ to zero in the simulation. For the widths of the Higgs bosons $H$ and $h$, we use an estimated tree-level total width, consisting of the partial widths for $h\rightarrow VV^{\left(*\right)}$, $h\rightarrow ff$, and $H\rightarrow hh$ and an effective $h\rightarrow\gamma\gamma$ partial width [^5] The total width as the sum of these partial widths is calculated using the “AutoWidth” feature of MadGraph5_amc@nlo.. As the current model file does not include the gluonic decay mode automatically, we also sum this decay width by hand, where we use the full top quark mass dependence and perform the calculation at LO. All partial widths were calculated with MadGraph5_amc@nlo. Some example values are shown in Table 1.
@@ V.1 Illustration of interference Effects
For illustrative purposes, we show the results for two parameter points, namely the points $m_{H}=300\,{\,\rm GeV}$, $\sin\alpha=0.17$, and $\tan\beta=3.3$; and $m_{H}=600\,{\,\rm GeV}$, $\sin\alpha=0.17$, and $\tan\beta=1.6$. For these points, we simulated $10^{6}$ events for each subprocess, $P_{\text{tot}}$, $P_{\mathrm{S_{H}}}$ and $P_{\text{woH}}$. We investigate the interference effects for different observables directly by evaluating the difference between the coherent sum $P_{\text{tot}}$ of all contributions to $gg\to hh$ and the incoherent sum $P_{\mathrm{S_{H}}}+P_{\text{woH}}$. In particular, we analyse the $m_{hh}$, $p_{\mathrm{T}}$, and $\Delta\phi_{hh}$ distributions, where $\Delta\phi_{hh}$ is the angle between the final state Higgs bosons. Here, $p_{\mathrm{T}}$ is shown for both final state Higgs bosons. While the behaviour of the di-Higgs invariant mass is well known, and the corresponding interference effects have been discussed in detail in the literature (see e.g. Ref. DiMicco:2019ngk and references therein), other variables have not been investigated yet. However, all of these can in principle serve as input for both cut-based as well as machine learning based analyses. Therefore, it is imminent to equally be aware of interference effects in as many as possible differential distributions.
@@
As we will discuss in the following, $m_{hh}$ and $p_{\mathrm{T}}$ show interesting interference effects, as can be seen in Figure 3, which shows the number of total events expected using full Run-2 luminosity = $\int\mathcal{L}\,=\,139\,{\rm fb}^{-1}$. In contrast, $\Delta\phi_{hh}$ is not significantly affected for these parameters points, and we chose to defer the corresponding discussion to Appendix A. We observe the well-known peak-dip structure around $m_{hh}\,\sim\,m_{H}$, that has already been discussed in the literature. For $m_{hh}$ slightly larger than the mass of the heavy scalar, there is a negative interference that leads to a smaller number of events than expected from incoherently adding the different distributions. A similar effect appears for $m_{hh}$ slightly below $m_{H}$, although the interference effect has a different sign. These are therefore regions in the $m_{hh}$ distributions where interference effects can become important. The interference behaves similarly for the $p_{\mathrm{T}}$ distributions, with positive interference for $p_{\mathrm{T}}$ lower than the peak and negative interference for $p_{\mathrm{T}}$ higher than the peak.
@@
Figure 3: The upper plots show the invariant mass $m_{hh}$ and the lower plots show the transversal momenta $p_{\mathrm{T}}$ of both final state Higgs bosons for the two parameter points. The blue line corresponds to $N_{\text{tot}}$, the red line corresponds to $N_{\mathrm{S_{H}}}$, the orange line corresponds to $N_{\text{woH}}$ and the green line is the sum $N_{\mathrm{S_{H}}}+N_{\text{woH}}$ neglecting interference effects. The event numbers $N_{i}$ corresponding to process $P_{i}$ were simulated assuming full Run-2 luminosity. The bin size is $10\,{\,\rm GeV}$.
@@
Another important question is whether these effects are statistically significant. In Figure 4, we therefore show the relative difference for the event numbers for the full Run-2 luminosity. As a simple estimate, we only take statistical errors into account. We see that the interference effects are non-negligible over large regions of parameter space, and exceed the statistical uncertainty estimate we applied.
@@
Figure 4: These plots show the interference calculated from the difference between the blue line and the green line seen in Figure 3 as the green line and a statistical error calculated as the $\sqrt{N_{\text{tot}}}$ of the $P_{\text{tot}}$ process assuming full Run-2 luminosity. The upper plots show this for the invariant mass $m_{hh}$ and the lower plots show this for the transversal momentum $p_{\mathrm{T}}$ for the two parameter points.
@@ V.2 Parameter scan
To investigate the effects of interference between the resonant and non-resonant diagrams, we examine the cross-section corresponding to the interference term for a broad range of parameters. For this purpose, we perform a parameter scan for a range of the relevant parameters in the parameter volume of $[260\,{\,\rm GeV},800\,{\,\rm GeV}]\times[-0.24,0.24]\times[0.5,4.0]$ of the $\{{m_{H},\sin\alpha,\tan\beta\}}$ space where we scan $m_{H}$ in steps of $60$ GeV, $\sin\alpha$ is steps of $0.08$ and $\tan\beta$ in steps of $0.5$. The resulting parameter points are displayed in Table 2. We omitted the points with $\sin\alpha=0$ as they represent the SM case where the second Higgs decouples.
@@
Figure 5: The different bins on the x-axis show the investigated values of $m_{H}$. The points inside the bins are ordered by the corresponding $\tan\beta$ values from left to right. The y-axis shows $\sin\alpha$, where the SM case of $\sin\alpha=0$ is omitted. Circles represent allowed points and triangles represent the points that do not pass the scan constraints described in Section II. The maximal relative interference for allowed points is around $13\%$. The colours of the points indicate the size of $\left(\Delta\sigma\right)_{\text{abs}}$ and $\left(\Delta\sigma\right)_{\text{rel}}$ for the left and right plots, respectively.
We investigated the interference effects for events with invariant mass $m_{hh}$ above or below $m_{H}$. For this we used the pylhe pylhe python package to divide our generated events into those with higher and those with lower invariant mass than the respective $m_{H}$. Accordingly, we split the cross-sections into the lower invariant mass range, $\left(\Delta\sigma\right)_{\text{abs}}^{<}$, and for the higher invariant mass range, $\left(\Delta\sigma\right)_{\text{abs}}^{>}$, respectively, as well as their absolute sum
@@
and the corresponding relative interferences $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, $\left(\Delta\sigma\right)_{\text{rel}}^{>}$ and $\left(\Delta\sigma\right)_{\text{rel}}^{\sum}$, defined accordingly by normalisation to the total cross-section over the whole mass range, equivalent to Eq. (24). The results are displayed in Figure 6. As shown in Figures 3a and especially 3b, the interference seems to have a positive value for $m_{hh}<m_{H}$ and a negative value for $m_{hh}>m_{H}$ [^7] In other beyond the SM extensions, one can also find scenarios where this behaviour is reversed, see e.g. discussion in Ref. Arco:2022lai in the context of a two-Higgs-doublet model.. Therefore, although the total interference integrated over the whole mass range might be small in some scenarios, effects in differential distributions could still be large. We therefore also provide the corresponding information on the different mass regions discussed above as well as the absolute value of deviations between naive summing of the different contributions and the total process including interference. Figure 6a shows that a higher positive mixing angle enhances both $\left(\Delta\sigma\right)_{\text{abs}}^{<}$ and $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, see Figure 6b. Furthermore, Figure 6c shows that $\left(\Delta\sigma\right)_{\text{abs}}^{>}$ has the largest negative values for large mixing angles in the $m_{H}$ region around $380\,{\,\rm GeV}$, see also Figure 6d. Figure 6e and especially Figure 6f show that some points, which appeared to have negligible interference based on the total cross-sections ($\left(\Delta\sigma\right)_{\text{abs}}$ and $\left(\Delta\sigma\right)_{\text{rel}}$) displayed in Figures 5a and 5b, actually experience two interference effects of opposite sign, which partially cancel when the total cross-section is computed. In Section VII.2 we will introduce benchmark (BM) points that exhibit this feature (benchmarks BM2 and BM3).
@@
Figure 6: Interferences for invariant di-Higgs masses $m_{hh}$ above and below the heavy scalar mass $m_{H}$. Bins on the x-axis represent different $m_{H}$ values, ordered by increasing $\tan\beta$ within each bin. The y-axis shows $\sin\alpha$, excluding $\sin\alpha=0$ (SM case). Circles represent allowed points and triangles represent the points that do not pass the scan constraints described in Section II. The colours of the points indicate the size of $\left(\Delta\sigma\right)_{\text{abs}}^{<}$, $\left(\Delta\sigma\right)_{\text{rel}}^{<}$, $\left(\Delta\sigma\right)_{\text{abs}}^{>}$, $\left(\Delta\sigma\right)_{\text{rel}}^{>}$, $\left(\Delta\sigma\right)_{\text{abs}}^{\sum}$, and $\left(\Delta\sigma\right)_{\text{rel}}^{\sum}$ for the upper-left, upper-right, middle-left, middle-right, lower-left, and lower-right, plots, respectively.
@@
Figure 7: The different bins on the x-axis show the different values of $m_{H}$ investigated and inside the bins the points are ordered by the corresponding $\tan\beta$ values from left to right. The $\sin\alpha$ value is shown on the y-axis, where $\sin\alpha=0$ is left out, as it corresponds to the SM-case. Circles represent allowed points and triangles represent the points that do not pass the scan constraints described in Section II. The colours of the points indicate the size of $\kappa_{\lambda_{hhh}}$, $\lambda_{Hhh}$, $\Gamma_{h}$, and $\Gamma_{H}$ for the upper-left, upper-right, lower-left, and lower-right plots, respectively.